OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 2) = 4x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (igreen):

\(\dfrac{2}{5}(x - 2) = 4x\) Distribute 2/5 into the parenthesis: \(\dfrac{2}{5}x - \dfrac{4}{5} = 4x\) Subtract 2/5x to both sides, what's 4x - 2/5x?

OpenStudy (anonymous):

-2/9

OpenStudy (anonymous):

hottie can i get a medal pweez?

OpenStudy (anonymous):

@mbailey436